OpenStudy (anonymous):

Solve Graph & Use Substitution or elimination? 3y=x+6 Y+2=-x it says solve and use either elimination or substiution. for the first problem I divided all of the problems by 3. but I don't know if its right because the slope would be x/3?? Help! I got the same answer for substitution and on the calc (-3,1) but I couldn't graph it I kept getting (-1,-1)?

OpenStudy (texaschic101):

y = -x - 2 now sub this into the other equation 3y = x + 6 3(-x - 2) = x + 6 -3x - 6 = x + 6 -6 - 6 = x + 3x - 12 = 4x - 3 = x now sub -3 in for x y = - x - 2 y = -(-3) - 2 y = 3 - 2 y = 1 Your ordered pair is (-3,1)

OpenStudy (anonymous):

Thanks how would I graph this though? Like I got it in slope intercept form by: Y= x/3 +2 Y= -x-2. Is it right that slope= x/3 that seems weird to me. I got that same pair (-3,1) but not the graph? Thank you!

OpenStudy (texaschic101):

y = mx + b 3y = 1x + 6 --> x is the same as 1x y = 1/3x + 2 slope is 1/3 and y intercept is 2

OpenStudy (anonymous):

Ok thanks alot!!

OpenStudy (texaschic101):

your welcome..glad to help :)